OpenStudy (anonymous):

I have been working on this forever i cant seem to find the intervals! find the area of the region 2 y = 4 \sqrt{x},y = 5, 2 y + 2 x = 6

OpenStudy (dumbcow):

first solve each one for y y=5 y = 2/sqrt(x) y = 3-x 3-x = 5 --> x = -2 2/sqrt(x) = 5 --> x = 4/25 3-x = 2/sqrt(x) look at graph, notice they intersect close to 1 check 3-1 = 2/1 2=2 --> x=1 \[A = \int\limits_{-2}^{4/25}5-(3-x) dx + \int\limits_{4/25}^{1}\frac{2}{\sqrt{x}}-(3- x)dx\]